OpenStudy (anonymous):

A and B are two sets. Given n(A union b)=36, n(A intersect B)=4 and n(B-A)=3[n(A-B)]

hero (hero):

What are you supposed to do?

OpenStudy (anonymous):

find n(A) and n(B)

hero (hero):

What course are you taking?

OpenStudy (anonymous):

form 6.. in malaysia

hero (hero):

I've never heard of that

hero (hero):

Sheg, any ideas?

OpenStudy (anonymous):

is a pre-u course

hero (hero):

A pre-who what? Yeah, I'm lost

OpenStudy (anonymous):

pre-university

hero (hero):

I see. Well, unless you can tell me what specific topic you are studying in your course, I doubt if I will be able to help

OpenStudy (anonymous):

number and sets..

OpenStudy (anonymous):

sub chapter sets..

OpenStudy (anonymous):

yeah n(AUB) = n(A) +n(B) -n(AnB) 36 = n(A) + n(B) - 4 n(A) + n(B) = 40

OpenStudy (anonymous):

yes hero was trying to recoll the rules of set theory

hero (hero):

Oh, I'm an idiot. I have done this before.

hero (hero):

Good job Sheg

OpenStudy (anonymous):